1. Assume an goods and services market of an economy is characterized by the following equations:

C = 0.8 (Y - T) I = 800 -20r Y=C+I+G T = 1000 G = 1000

a) Consider for the moment the Keynesian Cross model. What will happen to the GDP if G increases by 200? What is the multiplier?

b) Keep considering the Keynesian Cross model. What will happen to the GDP if T increases by 200? What if both G and T increase by 200?
Answer #1

C: 0.8(4-1)-,I=800- 20€, 6=1000, T: 10.00 Multiflies & I 1-4 1-0.8 ²1 02 - 5 Ale AY 5 If G increases G inreased by 200, thenIf I increases by 200, New 7 = 1200 Ye 0.844-1200) + 800-2000) tloos Yeo-84-960 + 800 -20 + lood 0.24 = 820 Y = 8200 4100 oft